# Critical temperature of the classical $XY$ model via autoencoder latent space sampling

### v1

@@ II Autoencoders
Before we show two use cases for physical systems, we demonstrate the use of autoencoders for generating a coordinate system for the sphere. Consider the set of points $\{(x,y,z)\in\mathbb{R}^{3}|x^{2}+y^{2}+z^{2}=1\}$ embedded in $\mathbb{R}^{3}$. We know that this is a unit sphere which is two-dimensional, however, this definition requires three parameters. We could easily use spherical polar coordinates $(\theta,\phi)$ but we use an autoencoder to find a different coordinatization. This autoencoder, shown in Fig. 2, is designed with three inputs $(x,y,z)$ and three outputs $(x^{\prime},y^{\prime},z^{\prime})$. These represent the points on the unit sphere embedded in $\mathbb{R}^{3}$. The latent space has two nodes since we require a two-dimensional coordinatization of the sphere.
Figure 2: A schematic of an autoencoder designed to map an embedded unit sphere into $\mathbb{R}^{2}$.
The loss function was set to
$L=\frac{1}{2}\Big{(}(x-x^{\prime})^{2}+(y-y^{\prime})^{2}+(z-z^{\prime})^{2}\Big{)}.$ (5)
The autoencoder was trained by passing in randomly sampled points from the unit sphere. After training, the latent space, as depicted in Fig. 3 and Fig. 5, was analysed by passing in paths on the sphere and observing these paths in the latent space. This was used to deduce the mapping that was learnt by the autoencoder.
Figure 3: Several paths around the sphere were used to analyse the latent space. These were a circle around the sphere in the $xy$ (red) plane, the $xz$ (blue) plane and the $yz$ plane (green). The intersections of these curves in the latent space are shown as labelled points in black.
We see from this latent space that the autoencoder has flattened the unit sphere, as seen in Fig. 4, into two dimensions, as seen in Fig. 5. The regions bounded by three different curves represents the octants of the original unit sphere.
Figure 4: The unit sphere in $\mathbb{R}^{3}$ with the first octant highlighted. Below the latent space with the region corresponding to the first octant on the unit sphere coloured in yellow.
Figure 5: The latent space with the region corresponding to the first octant on the unit sphere coloured in yellow.
@@ III The Model
Figure 6: Here we have examples of vortex (left) and anti-vortex (right) configurations on the lattice. In the continuum limit, these become the singular solutions (topological defects) mentioned above.
@@
As already stated, there is no spontaneous symmetry breaking at finite temperature, however, there is a transition between long-range correlations at low temperature and short range correlations at high temperature. Kosterlitz and Thouless[21] showed that at low temperatures the vortices occur in tightly bound pairs. As the temperature increases past a transition point, $T_{KT}/J\approx 0.893$, the pairs undergo deconfinement which results in a change in the order parameter from a power-law to exponential.
@@ IV Generating and sampling the latent space
Figure 7: The autoencoder was designed with a 60-dimensional latent space.
@@
where $D(x,y)$ is a disk of radius $\sigma$ centred at $(x,y)$ and $\sigma$ is a length on the order of the width of a vortex/anti-vortex. This field quantifies the average variation of $\theta(x,y)$ around a local neighbourhood of radius $\sigma$. $N(u,v)$ is a Gaussian centred at $(x,y)$. This term ensures that only field values close to $(x,y)$ are considered in the averaging process. The cosine term is analogous to the term in Eq. (7) and is used to remove the unwanted symmetry. This term also ensures that the field is bounded which will be important when implementing the autoencoder. An example of this $\theta(x,y$) field and its corresponding $A(x,y)$ field are illustrated in Fig. 8 and Fig. 9 respectively. Vortices and anti-vortices will have large field variations in their neighbourhood and so will result in a large field value. Regions with no vortices or anti-vortices will have little to no variation in the field. This will result in a very small field value. We can thus characterize the vortex/anti-vortex density by the magnitude of $A(x,y)$ across the extent of the field.
The autoencoder is then trained using these fields that are derived from $\theta(x,y)$ fields generated using standard MCMC methods. We generate a lattice where each site is given a random value $\theta(x,y)\in[0,2\pi]$. A single MCMC step involves choosing a random site, proposing a new site value and then accepting this change if the energy decreases or if a uniform random variable is less than $\exp(-\Delta E/T)$. This is the Metropolis-Hastings algorithm that allows us to sample from the required Boltzmann distribution. This is done for each temperature value.
Figure 8: This $\theta(x,y)$ field was generated using MCMC at a temperature $T=0.6$. The field is given by the vector field $(\cos\theta,\sin\theta)$. The background colours show the local energy at each point. A vortex can be seen in the lower left corner.
Figure 9: This $A(x,y)$ field was generated from the $\theta(x,y)$ field in Fig. 8. A low field value corresponds to little variation in the original field and a large field value corresponds to large variation.
@@
Once trained, the latent space was analysed by passing in $A(x,y)$ fields and generating histograms from the produced latent space values. These histograms were then used to determine the mean and standard deviation of the latent space variables. New fields can be generated by passing in a sampled latent space vectors, $z=(z_{1},z_{2},z_{3},...,z_{60})$, into the decoder portion of the autoencoder. Each $z_{i}$ is sampled from the respective Gaussian with mean $\mu_{i}$ and standard deviation $\sigma_{i}$. The particular Gaussian distribution for $z_{1}$ is shown in Fig. 10 below.
@@
Figure 10: Histogram of $z_{1}$ values from the latent space point $(z_{1},...,z_{60})$. Similar distributions can be obtained for the other dimensions of the latent space.
This sampling process is very computationally inexpensive compared to the standard MCMC algorithm. These sampled fields can then be used to calculate the thermodynamic average of the vortex density at a given temperature. We cannot directly count the number of vortices so instead we use the average,
@@
as a proxy which is plotted in Fig. 11. $A(x,y)$ is analogous to the energy at and around the point $(x,y)$. If each “vortex” has energy $\epsilon$ then $\frac{1}{\epsilon}\braket{A}=n$ gives the average number of unbound votices over the extent of the field. This will not be exact but will allow us to see certain trends that give an approximation to the transition temperature.
Figure 11: $\braket{A}$ is zero up to the transition temperature after which it increases rapidly. This transition temperature agrees well with that found in the literature, $\frac{k_{b}T_{c}}{J}=0.8935$ [26].
@@ V Conclusion
It was shown that an autoencoder can be a useful tool in reducing a given configuration space to a lower dimensional latent space that may be much easier to sample from. In this case, it was found that the latent space could be sampled using various Gaussian distributions. This latent space was sampled to calculate the thermal average of the vortex density and from this, one could determine the critical temperature ($T_{KT}$) at which this vortex density becomes non-zero. This method of latent space sampling is very general and so can be applied to many systems. The only requirement is that the systems need to contain a large enough amount of correlation between its constituents in order for the autoencoder to learn how to compress it with little loss. A large amount of correlation results in a large amount of redundant information can be removed during the compression stage of the autoencoder. With this in mind, one can extend this method to systems of large particles with solid-liquid phase transitions or systems with topological phase transitions [27] since one does not need any order parameters during the training process. Systems with very little to no correlations (like an ideal gas) cannot be mapped to a lower dimensional space since the knowledge of the behaviour of part of the system gives no information on the behaviour of another part. The specification of the system cannot be reduced.

### v2

@@ III The Model
Figure 2: Here we have examples of vortex (middle right) and anti-vortex (upper left) configurations on the lattice. In the continuum limit, these become the singular solutions (topological defects) mentioned above. The background colour shows the $\theta(x,y)$ field.
@@
As already stated, there is no spontaneous symmetry breaking at finite temperature, however, there is a transition between long-range correlations at low temperature and short range correlations at high temperature. Kosterlitz and Thouless[21] showed that at low temperatures the vortices occur in tightly bound pairs. As the temperature increases past a transition point, $k_{B}T_{KT}/J\approx 0.893$, the pairs undergo deconfinement which results in a change in the order parameter from a power-law to exponential.
@@ IV Generating and sampling the latent space
Figure 3: The autoencoder was designed with a 60-dimensional latent space. The architecture and training details are elaborated on in the appendix.
@@
where $D(x,y)$ is a disk of radius $\sigma$ centred at $(x,y)$ and $\sigma$ is a length on the order of the width of a vortex/anti-vortex. This field quantifies the average variation of $\theta(x,y)$ around a local neighbourhood of radius $\sigma$.
@@
$N(u,v)$ is a Gaussian centred at $(x,y)$. This term ensures that only field values close to $(x,y)$ are considered in the averaging process. The cosine term is analogous to the term in Eq. (6) and is used to remove the unwanted symmetry. This term also ensures that the field is bounded which will be important when implementing the autoencoder. An example of this $\theta(x,y$) field and its corresponding $A(x,y)$ field are illustrated in Fig. 4. Vortices and anti-vortices will have large field variations in their neighbourhood and so will result in a large field value. Regions with no vortices or anti-vortices will have little to no variation in the field. This will result in a very small field value. We can thus characterize the vortex/anti-vortex density by the magnitude of $A(x,y)$ across the extent of the field. The autoencoder is then trained using these fields that are derived from $\theta(x,y)$ fields generated using a standard MCMC method which is explained in the appendix. Other configuration generation methods such as those by Swendsen and Wolff [26, 27] can be used to generate the training data. However, as long as a sufficient amount of training samples is generated, the precious method of generation is not important.
Figure 4: This $\theta(x,y)$ field was generated using MCMC at $T=0.9$. The background is the auxiliary $A(x,y)$ field calculated using Eq. 13.
@@
Once trained, the latent space was analysed by passing in $A(x,y)$ fields and generating histograms from the produced latent space values for each dimension of the latent space. These histograms were then used to determine the mean and standard deviation of the distribution for each latent space dimension. New fields can be generated by passing in a sampled latent space vectors, $z=(z_{1},z_{2},z_{3},...,z_{60})$, into the decoder portion of the autoencoder. Each $z_{i}$ is sampled from the respective Gaussian with mean $\mu_{i}$ and standard deviation $\sigma_{i}$. The particular Gaussian distribution $N(\mu_{1}=0.14,\sigma_{1}=0.17)$ for the latent space dimension $z_{1}$ is shown in Fig. 5 below.
@@
Figure 5: Histogram of $z_{1}$ values sampled from the first latent space dimension. Similar distributions can be obtained for the other dimensions of the latent space as explained above.
This method is similar to, but technically different from, a variational autoencoder [29]. Through the training process, an autoencoder learns to map a high dimensional vector to a lower dimensional latent space. The exact process is not known explicitly but is manifested in the learned parameters. On the other hand, a variational autoencoder first maps the input to a mean vector and a standard deviation vector of a predefined distribution. This distribution is then sampled to produce the latent space vector. The generation of these latent space distributions can be done using either method however it should be noted that you do not need to predefine a distribution over the latent space before training when using an autoencoder. The latent space sampling process is very computationally inexpensive compared to the standard MCMC algorithm. These sampled fields can then be used to calculate the thermodynamic average of the vortex density at a given temperature. We cannot directly count the number of vortices so instead we use the average,
@@
as a proxy. $A(x,y)$ is analogous to the energy at and around the point $(x,y)$. If each “vortex” has energy $\epsilon$ then $\frac{1}{\epsilon}\braket{A}=n$ gives the average number of unbound votices over the extent of the field. Ideally, we would expect the function $\braket{A}$ to be
$\braket{A}=\Big{\{}\begin{array}[]{lr}0&\text{if }T<T_{c}\\ a(T-T_{c})&\text{if }T\geq T_{c}\end{array}$ (15)
where we have vortex unbinding above the critical temperature. The number of vortices, and hence vortex density, will then grow linearly with temperature. In practice, finite size effects and finite sampling effects will result in an approximate form of $\braket{A}$ as shown in the figure below.
Figure 6: We calculate $\braket{A}$ as a function of $T$ over the temperature range $20<T<$ 2. This was done for field sizes $N=25,50,75$ and 100. The straight line in black is a line of best fit.
Note the approximate fitting to Eq. 15. We can use the intercept point of the line of best fit as an approximation for $T_{c}$. The results are summarized in the table below.
$N$ $T_{c}$ Estimate Error $\Delta T$
25 0.872833 0.020067
50 0.872213 0.020687
75 0.871779 0.021121
100 0.872756 0.02144
Table 1: Table of critical temperature estimates and associated errors for varying system sizes.
@@ V Conclusion
It was shown that an autoencoder can be a useful tool in reducing a given configuration space to a lower dimensional latent space that may be much easier to sample from. In this case, it was found that the latent space could be sampled using various Gaussian distributions. This latent space was sampled to calculate the thermal average of the vortex density and from this, one could determine the critical temperature ($T_{KT}$) at which this vortex density becomes non-zero. This method of latent space sampling is very general and so can be applied to many systems. The only requirement is that the systems need to contain a large enough amount of correlation between its constituents in order for the autoencoder to learn how to compress it with little loss. A large amount of correlation results in a large amount of redundant information can be removed during the compression stage of the autoencoder. With this in mind, one can extend this method to systems of large particles with solid-liquid phase transitions or systems with topological phase transitions [30] since one does not need any order parameters during the training process. Systems with very little to no correlations (like an ideal gas) cannot be mapped to a lower dimensional space since the knowledge of the behaviour of part of the system gives no information on the behaviour of another part, ie. the specification of the system cannot be reduced. Future work for this method includes extensions to other learnable features such as magnetization where a magnetization density vector field $M(x,y)$ is derived from the $\theta(x,y)$ field instead of the auxiliary $A(x,y)$ field as done in this paper. Other work includes extensions to the generalized XY model which includes fractional vorices. In that case, a new vortex counting method needs to be implemented.
VI Appendix
Pseudo code for MCMC generation
Ψ#Generate a field of size N x N at temp. T
ΨTWO_PI = 6.283185
ΨN = 100
ΨT = 0.6
ΨSTEPS = 1000 #Number of MCMC steps
Ψfield[N][N]  #Field configuration
Ψ
Ψ#Initialize the field randomly
Ψ#rand() returns a uniform random float on (0,1)
Ψfor i = 0 to N
ΨΨfor j = 0 to N
ΨΨΨfield[i][j] = rand() * TWO_PI
Ψ
Ψfor step = 0 to STEPS
ΨΨ#Choose a random site
ΨΨi = rand() * N; j = rand() * N
ΨΨE_prev = -cos(field[i][j]-field[i+1][j])
ΨΨ         -cos(field[i][j]-field[i-1][j])
ΨΨ         -cos(field[i][j]-field[i][j+1])
ΨΨ         -cos(field[i][j]-field[i][j-1])
ΨΨ#Propose a small change to the field
ΨΨdelta = rand() * TWO_PI * 0.1
ΨΨfield[i][j] += delta
ΨΨE_new = -cos(field[i][j]-field[i+1][j])
ΨΨ        -cos(field[i][j]-field[i-1][j])
ΨΨ        -cos(field[i][j]-field[i][j+1])
ΨΨ        -cos(field[i][j]-field[i][j-1])
ΨΨdE = E_new - E_prev
ΨΨ#Accept or Reject using Metropolis-Hastings
ΨΨp = min(1.0, exp(-dE/T))
ΨΨif rand() < p #accept
ΨΨelse #reject
ΨΨΨfield[i][j] -= delta
The $\theta(x,y)$ and $A(x,y)$ fields were discretized into lattices of sizes N = 25, 50, 75 and 100. The size of the autoencoder input layer for each lattice is $N^{2}$. The size of each subsequent layer is $\frac{3}{4}$ of the previous layer in order to create the required bottleneck for the auto encoder. The size of the latent space was chosen to be 60. This choice was a good enough compromise between computational cost and reconstruction detail. Quadratic cost (MSE) was chosen as the cost function for its simplicity. The training was done for 5000 epochs with a batch size of 100 and a learning rate of 0.001. The sigmoid activation function was used for all layers. Standard MCMC methods were then used to generate the training data. For each temperature value, 10000 samples were generated and used as training data. The full code (with comments) for this paper can be found on github.com/BrandonWillnecker
Acknowledgements.
We would like to acknowledge the insightful comments and suggestions made by the referees.